Parth (parthkohli):

You have two empty bags and 100 bread rings. The goal is to put all your rings into the bags in such a way that one of the bags will contain exactly twice as many rings then the other. You can't eat, break and lose the rings. All of them should go into the bags.

OpenStudy (asnaseer):

is the fact that they are "rings" significant?

OpenStudy (asnaseer):

oh wait - I think I have it!

OpenStudy (asnaseer):

but I'll wait to see if others can find this - nice problem :)

OpenStudy (asnaseer):

the solution I have involves the bread rings

OpenStudy (anonymous):

\[A + B = 100 \] \[A+B=100 = 2B + B = 3B = 100 B=\frac{ 100 }{ 3 }\] But no this can't be.....we would have to break rings for this...

OpenStudy (anonymous):

Well I think I have a solution.What if we split the number of rings and put them in each bag and ten put one bag into another? :) In this way the condition is satisfied.Anyway it could've been simpler if the number of rings were 99.

OpenStudy (asnaseer):

yes - that is the solution I had as well

OpenStudy (anonymous):

@asnaseer Cheers!

Parth (parthkohli):

@kartiksriramk has the correct solution :)